City Taxi Service purchased a new auto to use as a taxi on January 1, Year 1, for $22,200. In addition, City paid sales tax and title fees of $1,440 for the vehicle. The taxi is expected to have a five-year life and a salvage value of $5,990.

Required

a. Using the straight-line method, compute the depreciation expense for Year 1 and Year 2.
b & c. Assume that the taxi was sold on January 1, Year 3, for $19,067. Prepare the general journal entries to record the Year 1 depreciation and sale of the taxi in Year 3.
Answer #1

Total cost = 22200+1440 = 23640

a) Depreciation expense

Year 1 Year 2
Depreciation expense (23640-5990/5) = 3530 3530

b) Journal entry

No Account and explanation Debit Credit
b No entry
c Cash 19067
Accumulated depreciation-Taxi (3530*2) 7060
Gain on Sale of taxi 2487
Taxi 23640